# A Homothetic Gauge Theory

### v1

@@ Abstract
While the principle of local gauge invariance forms the bedrock of modern physics, the symmetry of scale invariance has traditionally been treated on a different footing. This paper develops a homothetic gauge theory (HGT) that elevates local scale transformations to a genuine gauge symmetry, placing them within the same geometric framework as conventional interactions. We introduce a novel mathematical structure based on an extended vector bundle that linearizes the action of local scaling, allowing for the definition of a homothetic connection and a corresponding gauge-covariant field strength. When this formalism is applied to classical U(1) electromagnetism, the theory yields a generalized field-strength tensor that naturally incorporates both the standard Maxwell field and a new scalar partner field governing the local scale. The resulting dynamics, including modified conservation laws and wave equations, are shown to be fully consistent. As a significant physical application, this framework provides a compelling resolution to the classical problem of the infinite self-energy of a point charge. By modeling the source as a penalized Dirichlet boundary condition on a shrinking interface, we construct a non-singular electrostatic potential, thereby removing the divergence in a natural, geometric manner without introducing ad-hoc regularization schemes. The emergence of penalty terms due to essentially geometric transformations promises a physically meaningful theory, and encourages to re-examining the role of scale invariance in other areas of classical and quantum field theory.
Keywords: homothetic gauge theory, scale invariance, conformal symmetry, gauge field, electromagnetic self-energy, geometric field theory, vector bundles, differential geometry
classification 11.15.-q, 11.30.-j, 02.40.-k, 03.50.De, 11.10.Ef
@@ 1 Introduction
An Operational Roadmap of the Paper: We begin by interpreting conformal geometry as a special case of a broader homothetic framework—specifically, as a homothetic geometry with a zero homothety center. This perspective allows us to elevate local scale transformations to a general gauge symmetry. To facilitate this, we introduce a novel mathematical structure based on doubled tangent and cotangent bundles, which linearizes the action of the homothetic group. Upon this foundation, we construct a set of gauge-covariant differential operators, leading to a generalized de Rham complex and an associated Hodge decomposition theorem, which ensures the mathematical consistency of the resulting field theories. As a primary application, we apply this formalism to classical electromagnetism, yielding a set of homothetically covariant Maxwell equations. A distinctive feature of these equations is the emergence of interaction terms that couple the electromagnetic field to the scalar gauge field governing the local scale. These terms persist even under simplifying assumptions and function as geometric penalty terms. This structure allows for the imposition of boundary conditions directly within the field equations, thereby reformulating the theory in a sourceless, singularity-free manner. By imposing a generalized Lorenz gauge condition, we derive homothetic wave and Laplace equations that naturally encode boundary data. This framework is shown to be equivalent to solving classical Dirichlet or Neumann problems via a volumetric penalization method. Finally, we leverage this result to address the classical problem of the infinite self-energy of a point charge. By modeling the source as a penalized Dirichlet boundary condition on a shrinking interface, we construct a non-singular electrostatic potential, resolving the divergence in a manner that emerges naturally from the geometry.
@@
This paper is structured as follows. Section 2 introduces the block–homothetic gauge group and the doubled-bundle construction that renders local scale transformations linear. Section 3 embeds this algebraic setup in a generalized Weyl-integrable geometry, deriving the compatible connection, curvature, and homothetic cohomology. In Section 4 we develop the homothetic de Rham complex and prove an extended Hodge decomposition that underpins the subsequent field theory. Section 5 specializes the formalism to electromagnetism, obtaining homothetically covariant Maxwell equations and a modified Lorentz force. Section 6 then derives the associated conservation law together with the homothetic wave and Laplace equations, while Section 7 analyses their Cauchy, Dirichlet and Neumann boundary-value problems. These results are applied in Section 8 to construct a non-singular electrostatic potential for a point charge. Section 9 summarises the main findings. For the sake of continuity, some merely technical materials are gathered in four appendices at the end of the paper.
@@ 2 First setup: Block-homothetic Transformations
:=$\mathcal{H}:=\left\{\Lambda[\lambda]\mid\lambda\in C^{\infty}(M,\mathbb{R})\right\},\qquad\Lambda[\lambda](x):=\begin{pmatrix}e^{-\lambda(x)}&1-e^{-\lambda(x)}\\ 0&1\end{pmatrix},$ (3)
@@
Let $\bar{\bf X}:=\begin{pmatrix}\bar{X}\\ \bar{X}^{\flat}\end{pmatrix}$ and $\hat{\bf X}:=\begin{pmatrix}\hat{X}\\ \hat{X}^{\flat}\end{pmatrix}$. Then:
$\hat{\bf X}=\begin{pmatrix}\Lambda&0\\ 0&\Lambda\end{pmatrix}\begin{pmatrix}\bar{X}\\ \bar{X}^{\flat}\end{pmatrix}={\bf H}\cdot\bar{\bf X}.$ (7)
@@
¯$\bar{\flat}_{x}:T_{x}M\oplus T_{x}M\to T_{x}^{*}M\oplus T_{x}^{*}M,\qquad\bar{X}\mapsto\bar{X}^{\bar{\flat}}:=(\eta(X),\eta(X_{d}))=(X^{\flat},X_{d}^{\flat}).$ (8)
@@ 3 A Generalized Weyl Integrable Geometry
On the extended space $\bar{\bf X}$, a symmetry manifests itself: the transformation $\bar{\bf X}\to\hat{\bf X}={\bf H}\bar{\bf X}$ preserves the musical isomorphism $\bar{\flat}$. In fact, one can verify that $\bar{\flat}$ and $\Lambda$ commute:
@@
However, this does not imply that the lengths and angles of vectors are preserved under the transformation $\bar{\bf X}\to\hat{\bf X}$—only that they remain unchanged within a local Lorentz frame. To make this distinction precise, we define the effective (generalized Weyl integrable) metric $\tilde{\eta}$ for any $X,Y\in TM$ via:
$\tilde{\eta}(X,Y):=\eta(\tilde{X},\tilde{Y}),$ (12)
@@
This metric generalizes the usual conformal structure of Weyl integrable geometry by incorporating an explicit offset vector $X_{d}\in TM$. For $\lambda=0$, the effective metric reduces to $\tilde{\eta}=\eta$, and for $\lambda\to\infty$, the metric degenerates:
@@
independently of $X$ and $Y$. The structure encoded in Eq. (13) thus reflects a homothetic-invariant geometry, one that naturally combines a local scaling $\lambda(x)$ with a fiberwise translation defined by $X_{d}(x)$, consistent with the action of group $\mathcal{H}$.
While this offset structure is geometrically natural and consistent with our next cohomological framework, it is hardly interpretable physically. The general form of the metric in Eq. (13), with $X_{d}\neq 0$, represents the full mathematical structure that arises from the linearization of the homothetic map. It is the most direct consequence of our initial setup. However, to ensure that this effective metric $\tilde{\eta}$ can be interpreted as a physical spacetime metric, it must satisfy certain fundamental axioms. In particular, the presence of $X_{d}$ in $\tilde{\eta}$ means that even the zero vector $X=0$ can acquire a nontrivial norm—a violation of the identity of indiscernibles and an obstruction to interpreting $\tilde{\eta}$ as a standard spacetime metric. Therefore, to recover a physical spacetime geometry that supports unambiguous geodesic separation, we impose a physicality condition by restricting the general geometric structure to the subspace where the offset vector vanishes. This is achievable without altering the foundational structure by choosing:
@@
However, we see this conformal geometry as a Weyl integrable geometry [28] because we want to use it in generalizing the cohomology. Recall that a Weyl integrable geometry is defined by the triple $(M,[\eta],\phi)$, where $[\eta]$ is the conformal class of Lorentzian metrics and $\phi:=d\lambda$ is a globally defined, closed 1-form satisfying
@@
This non-metricity condition defines a Weyl-compatible connection that preserves null directions but allows local rescaling governed by $\phi$. In our formulation, both the conformal metric and the 1-form $\phi:=d\lambda$ arise naturally from the homothetic structure, with $\lambda\in C^{\infty}(M)$ arbitrary and smooth. Thus, the effective geometry obtained from Eq. (16) is not merely analogous to Weyl integrable geometry—it is precisely one, derived from a block-homothetic framework.
About the Physical Interpretation of $\lambda$
In the absence of any other valid physical interpretation, we posit that the physical interpretation of this conformal frame is that of a spacetime with a variable speed of light (VSL). Establishing a relation between the conformal transformation and the variable speed of light is almost trivial, for example, one can see the following.
Consider the normalization of the four-velocity $u^{\mu}$ for a massive particle. In the physical frame described by $\tilde{\eta}_{\mu\nu}$, the squared norm is given by $\tilde{\eta}_{\mu\nu}u^{\mu}u^{\nu}$. To maintain a canonical structure where this norm is identified with the negative square of the local speed of light, we must define a position-dependent speed of light $c(x)$ such that:
$\tilde{\eta}_{\mu\nu}u^{\mu}u^{\nu}=-c(x)^{2}.$ (18)
Comparing with $\eta_{\mu\nu}u^{\mu}u^{\nu}=-1$, this naturally leads to the identification:
$c(x)=e^{-\lambda(x)}.$ (19)
This VSL interpretation is further substantiated in the limit where $\lambda(x)\to\infty$. In this case, $c(x)\to 0$ and the metric components $\tilde{\eta}_{\mu\nu}\to 0$. The collapse of all spacetime intervals signifies a degeneration of the causal structure where propagation is impossible, consistent with a vanishing speed of light.
Crucially, the existence of a relativistic kinematic framework, regardless of the local value of $c(x)$, allows for the definition of its two fundamental degenerate limits. At any spacetime point $x$, the local physics is governed by a Lorentz algebra characterized by the finite, non-zero constant $c(x)$. The Galilean and Carrollian kinematics can be understood as distinct Inönü–Wigner contractions [17] of this local Poincaré algebra, corresponding to the limits where the structure constant $c(x)$ is taken to infinity or zero, respectively.
  • The Galilean Limit ($c(x_{0})\to\infty$): This contraction yields the Galilean algebra $\mathfrak{g}(1,3)$. Physically, it corresponds to the causal cone opening up completely, allowing for absolute simultaneity and instantaneous propagation of signals.
  • The Carrollian Limit ($c(x_{0})\to 0$): This contraction yields the Carrollian algebra $\mathfrak{c}(1,3)$. Physically, it corresponds to the causal cone collapsing to a one-dimensional line, forbidding motion and rendering causality ultra-local.
This establishes that the Galilean and Carrollian frameworks are universally definable as conceptual limits from any local relativistic standpoint. The role of the field $\lambda(x)$ is then to dictate the physical character of a given spacetime region. While an observer can always conceive of both limits, a region where $\lambda(x)\to\infty$ will see its physical dynamics—its particle propagation and causal structure—asymptotically approach the behavior described by the idealized Carrollian limit [22].
@@ 4 Homothetic Hodge Decomposition
In this section, we study the behavior of differential forms in the homothetically transformed geometry introduced previously. Rather than focusing on traditional invariance properties, we consider a class of homothetically transformed differential forms, which lead naturally to a generalized de Rham complex and cohomology, and a generalized Hodge decomposition, adapted to the structure. We begin with the case of 1-forms, establishing both their physical interpretation and mathematical formulation. Other $k\neq 1$ ranks will be addressed in subsequent subsections.
4.1 The Homothetically Transformed 1-Forms
Finally, the present paper intends to analyze electromagnetism in which the (Ehresmann) connection, on which we can put physically meaningful constraints, is ${\mathfrak{u}}(1)$-valued 1-forms $A=A_{\nu}dx^{\nu}\in\Omega^{1}(M)$. Therefore, we begin here with 1-forms, albeit in a more general setting which will be suppressed to the potential gauge fields $A$ in §$\S$ 5.
In the conventional formulation of conformal geometry, the spacetime metric rescaling is a postulation occurring independently of the definition of other fields. But here, the conformal rescaling $\tilde{\eta}\to e^{-2\lambda}\eta$ is a consequence of a more fundamental transformation acting on the tangent and cotangent bundles. Since a 1-form $\alpha$ is a section of this very cotangent bundle, it is not only legitimate but more consistent within our framework to posit that it transforms accordingly. We therefore define the action of the homothetic group $\mathcal{H}$ on the space of extended 1-forms. For a general $\alpha\in\Omega^{1}(M)$, its transformed counterpart $\tilde{\alpha}$ is given by the homothetic transformation:
@@ 4.1 The Homothetically Transformed 1-Forms
where $\alpha_{d}\in\Omega^{1}(M)$ is a fixed, globally defined 1-form that serves as the homothety center, in direct analogy to the covector $X^{\flat}_{d}$ in the transformation of cotangent spaces. Note that when $\alpha_{d}=0$, this reduces to the familiar Weyl-type scaling $\tilde{\alpha}=e^{-W\lambda(x)}\alpha$, with the Weyl weight $W=1$. But in contrast to $X_{d}$, here we will keep $\alpha_{d}$ non-zero in general.
Before continuing, it is important to note that in this formulation, a homothetically invariant 1-form is obtained by setting:
@@
that is, $\tilde{\alpha}$ for $\lambda(x)=0$ equals $\tilde{\alpha}$ for $\lambda(x)\to\infty$, which means that changing $\lambda(x)$ does not affect $\alpha$. This can also be derived from definition (20):
@@
As will be seen, even for homothetically invariant forms (such as electromagnetic gauge potentials, as discussed in §5), the general homothetic formulation reveals new and interesting features, as it localizes the values of $\alpha$.
The physical motivation and mathematical justification for the homothetic transformation (20) are now in order.
Physical motivation. Consider the limiting case $\lambda(x)\to\infty$, where the conformal factor goes to zero, $e^{-2\lambda(x)}\to 0$, and hence the associated volume form goes to zero as well $(d\text{vol}_{\tilde{\eta}})_{x}\to 0$. Now, in order to have a theory, being valid in all scales, the functions on $(d\text{vol}_{\tilde{\eta}})_{x}$ must converges to a single well-defined value $(\alpha_{d})_{x}$. Note that such a homothetic behavior may define a kind of homogenization, that is, the transformation in Eq. (20) ensures that for any initial 1-form $\alpha$, its transform $\tilde{\alpha}$ approaches the fixed form $\alpha_{d}$:
@@
This principle thus endows the asymptotic field $\alpha_{d}$ with a clear physical interpretation: it is the universal 1-form to which all others are driven as the local spacetime structure degenerates. The existence of such a non-trivial asymptotic field is a profound consequence of the homothetic framework.
Mathematical formalism. This intuition can be rigorously framed using contraction on the cotangent bundle. Define, at each point $x\in M$, a fiberwise homothetic map:
@@
Endow each fiber $T_{x}^{*}M$ with the norm $\|\cdot\|_{x}$ induced by the original Minkowski metric $\eta$. Then for $\lambda(x)>0$, ${\mathbf{T}}_{\lambda(x)}$ is a strict contraction with Lipschitz constant $e^{-\lambda(x)}<1$. By the Banach fixed-point theorem, ${\mathbf{T}}_{\lambda(x)}$ has a unique fixed point in each fiber, which is precisely $\alpha_{d}$. Moreover, as $\lambda(x)\to\infty$, ${\mathbf{T}}_{\lambda(x)}$ converges to a projection map that sends every 1-form to $\alpha_{d}$:
$\lim_{\lambda(x)\to\infty}{\mathbf{T}}_{\lambda(x)}(\alpha)=\alpha_{d},\quad\forall\alpha\in T_{x}^{*}M.$ (24)
This construction defines a smooth family of fiberwise contractions across $M$, forming a pointwise contraction semigroup. In the limit $\lambda(x)\to\infty$, $\tilde{\alpha}$ thus enforces homogenization of all 1-forms to a single invariant form.
Now, as we mentioned before (c.f., Eqns. (9) and (10)), the above homothetic transformation can be linearized via the embedding $\Omega^{k}(M)\hookrightarrow\Omega^{k}(M)\oplus\Omega^{k}(M)$, that is:
$\hat{\alpha}:=\Lambda\cdot\bar{\alpha}=\begin{pmatrix}\tilde{\alpha}\\ \alpha_{d}\end{pmatrix}=\begin{pmatrix}e^{-\lambda(x)}\alpha+\left(1-e^{-\lambda(x)}\right)\alpha_{d}\\ \alpha_{d}\end{pmatrix}.$ (25)
The implications of this contraction will become central in the upcoming discussion of generalized Hodge theory.
What happen to the other differential forms is the question that will be addressed in the next subsection, after developing some mathematical tools we need for their analyses.

### v2

@@ Abstract
While the principle of local gauge invariance forms the bedrock of modern physics, the symmetry of scale invariance has traditionally been treated on a different footing. This paper develops a homothetic gauge theory (HGT) that promotes local scale transformations to a genuine gauge symmetry, placing them within the same geometric framework as conventional interactions. We introduce a novel mathematical structure based on an extended vector bundle that linearizes the action of local scaling, allowing for the definition of a homothetic connection and a corresponding gauge-covariant field strength. A central mathematical innovation of this theory is the development of a homothetic Hodge–de Rham complex and cohomology, wherein the standard differential and codifferential operators are deformed in a scale-covariant manner. When this formalism is applied to classical electromagnetism, as a U(1) gauge theory, the generalized homothetic Maxwell equations are obtained, and a direct connection is established with the spontaneous symmetry breaking. As a significant physical application, this framework provides a compelling resolution to the classical problem of the infinite self-energy of a point charge. By modeling the source as a penalized Dirichlet boundary condition, we construct a non-singular electrostatic potential.
Keywords: gauge theory, scale invariance, local symmetries, vector bundles, field theory, Weyl integrable geometry, Hodge–de Rham cohomology, spontaneous symmetry breaking, Maxwell equations, potential theory
classification
@@ 1 Introduction
A feature of HGT is that it clarifies the usual split between symmetry of the equations and symmetry of their solutions [10, 25]. Instead of purely multiplicative Weyl rescalings, we employ an affine action generated by the dilatation one-form $b=d\lambda$ with offset fields $\phi_{d}$. The field equations remain homothetically covariant, while physically meaningful scales enter through initial/boundary data and gauge choices. In this manner, “spontaneous” breaking of local scale invariance arises from selecting a background/gauge (fixing $\lambda$) and offsets $\phi_{d}$ that act as order parameters, not from a Higgs-type potential. The equations keep homothetic covariance; particular solution classes pick a preferred scale via the imposed data (see $\S$ 7) [33, 25].
The generalized Maxwell sector (Secs. 5.1 &\& 7.1) makes the above issue concrete: the homothetic Lorenz gauge and associated wave/Laplace operators support a penalization that enforces Dirichlet data on a shrinking interface. In the singular limit one recovers the exterior electrovac and avoids the point-charge self-energy divergence; the emergent scale is set by the penalization profile and choice of $\lambda$ (Sec. 8.2), so symmetry is preserved in the equations but broken by solution selection.
An Operational Roadmap of the Paper
@@
We begin by interpreting conformal geometry as a special case of a broader homothetic framework—specifically, as a homothetic geometry with a zero homothety center. This perspective allows us to elevate local scale transformations to a general gauge symmetry. To facilitate this, we introduce a novel mathematical structure based on doubled tangent and cotangent bundles, which linearizes the action of the homothetic group. Upon this foundation, we construct a set of gauge-covariant differential operators, leading to a generalized de Rham complex and an associated Hodge decomposition theorem, which ensures the mathematical consistency of the resulting field theories. As a primary application, we apply this formalism to classical electromagnetism, yielding a set of homothetically covariant Maxwell equations. A distinctive feature of these equations is the emergence of interaction terms that couple the electromagnetic field to the scalar gauge field governing the local scale. These terms persist even under simplifying assumptions and function as geometric penalty terms. This structure allows for the imposition of boundary conditions directly within the field equations, thereby reformulating the theory in a sourceless, singularity-free manner. By imposing a generalized Lorenz gauge condition, we derive homothetic wave and Laplace equations that naturally encode boundary data. This framework is shown to be equivalent to solving classical Dirichlet or Neumann problems via a volumetric penalization method. Finally, we leverage this result to address the classical problem of the infinite self-energy of a point charge. By modeling the source as a penalized Dirichlet boundary condition on a shrinking interface, we construct a non-singular electrostatic potential, resolving the divergence in a manner that emerges naturally from the geometry.
@@
The paper is structured as follows. Section 2 introduces the block–homothetic gauge group and the doubled-bundle construction that renders local scale transformations linear. Section 3 embeds this algebraic setup in a generalized Weyl-integrable geometry, deriving the compatible connection and curvature. In Section 4 we develop the homothetic de Rham complex and prove an extended Hodge decomposition that underpins the subsequent field theory. Section 5 specializes the formalism to electromagnetism, obtaining homothetically covariant Maxwell equations and a modified Lorentz force. Section 6 then derives the associated wave and Laplace equations. Section 7 discusses the relation between HGT and spontaneous symmetry breaking in U(1)-symmetric theories. The physical and mathematical applications of the theory are provided in two related sections. Section 8.1 analyses the Cauchy, Dirichlet, and Neumann boundary-value problems; the discussions that are applied in Section 8.2 to construct a non-singular electrostatic potential for a point charge. Section 9 summarises the main findings. For the sake of continuity, some technical materials are gathered in five appendices at the end of the paper.
@@ 2 First setup: Block-homothetic Transformations
$\mathcal{H}:=\left\{\Lambda[\lambda]\mid\lambda\in C^{\infty}(M,\mathbb{R})\right\},\qquad\Lambda[\lambda](x):=\begin{pmatrix}e^{-\lambda(x)}&1-e^{-\lambda(x)}\\ 0&1\end{pmatrix}$, (3)
@@
Remark 1
One can write:
$\Lambda[\lambda(x)]\;=\;e^{-\lambda(x)T}=\;\begin{pmatrix}e^{-\lambda(x)}&1-e^{-\lambda(x)}\\[2.0pt] 0&1\end{pmatrix}$, (4)
where
$T\;=\;\begin{pmatrix}1&-1\\[2.0pt] 0&\,0\end{pmatrix}$. (5)
In the rest of the paper, everywhere that is not confusing, we enjoyed the shorthand writing $\bar{X}\mapsto e^{-\lambda(x)T}\,\bar{X}$, or for the fiberwise operations $X\mapsto e^{-\lambda}\,X$, for $X\in{\mathbb{R}}^{2}$.
@@
Let $\bar{\bar{X}}:=\begin{pmatrix}\bar{X}\\ \bar{X}^{\flat}\end{pmatrix}$ and $\hat{\hat{X}}:=\begin{pmatrix}\hat{X}\\ \hat{X}^{\flat}\end{pmatrix}$. Then:
$\hat{\hat{X}}=\begin{pmatrix}\Lambda&0\\ 0&\Lambda\end{pmatrix}\begin{pmatrix}\bar{X}\\ \bar{X}^{\flat}\end{pmatrix}={\bf H}\cdot\bar{\bar{X}}.$ (9)
@@
$\bar{\flat}_{x}:T_{x}M\oplus T_{x}M\to T_{x}^{*}M\oplus T_{x}^{*}M,\qquad\bar{X}\mapsto\bar{X}^{\bar{\flat}}:=(\eta(X),\eta(X_{d}))=(X^{\flat},X_{d}^{\flat}).$ (10)
@@ 3 A Generalized Weyl Integrable Geometry
On the extended space $\bar{\bar{X}}$, a symmetry manifests itself: the transformation $\bar{\bar{X}}\to\hat{\hat{X}}={\bf H}\bar{\bar{X}}$ preserves the musical isomorphism $\bar{\flat}$. In particular, it is due to the commutativity of $\bar{\flat}$ and $\Lambda$:
@@
However, this does not imply that the lengths and angles of vectors are preserved under the transformation $\bar{\bar{X}}\mapsto\hat{\hat{X}}$—only that they remain unchanged within a local Lorentz frame. To make this distinction precise, we define an affine symmetric 2-form $\tilde{\eta}:TM\times TM\to{\mathbb{R}}$ via:
[…]
@@
As one can see, this affine metric kernel reflects explicitly the existence of the offset vector $X_{d}$ and the scaling $\lambda(x)$. It is worth mentioning that for $\lambda=0$, it reduces to $\tilde{\eta}=\eta$, and for $\lambda\to\infty$, it degenerates to:
@@
independently of $X$ and $Y$. The structure encoded in Eq. (15) thus reflects the symmetry (13) that naturally combines a local scaling $\lambda(x)$ with a fiberwise translation defined by $X_{d}(x)$, consistent with the action of the group $\mathcal{H}$.
While this offset structure is geometrically natural and consistent with our next cohomological framework, however, the general 2-form (15), with $X_{d}\neq 0$, is hard to justify as a physical spacetime metric. It arises from the linearization of the homothetic map, and it is the most direct consequence of our initial setup. However, it is not a metric in the usual sense. In particular, the presence of $X_{d}$ in $\tilde{\eta}$ means that even the zero vector $X=0$ can acquire a nontrivial norm— violates the usual requirement that $\|0\|=0$. Therefore, to recover a physical spacetime geometry that supports unambiguous geodesic separation, we impose a physicality condition by restricting ourselves to the subspace where the offset (homothety center) vector vanishes:
@@
It should be stressed that although we finally set $X_{d}=0$, this is not just a complicated way to achieve the standard conformal geometry. In fact, we will always consider this geometry as a homothetic geometry with the zero homothety center. In particular, it manifests itself in the differential forms we will associate with it in the next section— see Eq. (20) and the discussions after that.
On the other hand, nothing in our formulation prevents us from admitting the conventional Weyl structure. Recall that [30] a Weyl integrable geometry is defined by the triple $(M,[\eta],\phi)$, where $\eta]$ is the conformal class of Lorentzian metrics and $\phi:=d\lambda$ is a globally defined, closed 1-form satisfying
@@
This non-metricity condition defines a Weyl-compatible connection that preserves null directions but allows local rescaling governed by $\phi$. However, in our formulation, both the conformal metric and the 1-form $\phi:=d\lambda$ arise naturally from the homothetic structure. In particular, the non-metricity 1-form $\phi=d\lambda$ appears automatically from the homothetic scaling parameter.
@@ 4 Homothetic Hodge Decomposition
In this section, the Hodge-de Rham theory is established for the homothetically transformed differential forms, defined on the above conformally transformed spacetime.
4.1 The Homothetically Transformed Differential Forms
Given the differential forms $\phi,\phi_{d}\in\Omega^{\bullet}(M)$, where $\phi_{d}$ is fixed, we define the homothetically transformed differential form $\tilde{\alpha}$ as:
@@ 4.1 The Homothetically Transformed Differential Forms
obtainable from the first row of $\hat{\alpha}=\Lambda\bar{\alpha}$. Associating such a transformation to the conformally transformed spacetime is mathematically legitimated (see appendix A and the discussion below). Moreover, the following special cases are noticeable:
  • For $\alpha_{d}=0$, it reduces to the familiar Weyl-type scaling $\tilde{\alpha}=e^{-W\lambda(x)}\alpha$, with the Weyl weight $W=1$. In contrast to $X_{d}$, here we will keep $\alpha_{d}\neq 0$ in general.
  • A homothetically invariant form (equivalent to $W=0$ in the classical Weyl geometry) is obtainable by setting:
@@
    that is, $\lambda(x)$ does not affect $\alpha$ because:
@@
    As will be seen, even for the homothetically invariant forms (such as electromagnetic gauge potentials, as discussed in $\S$ 5), the general homothetic formulation has still new and interesting things.
The physical motivation and mathematical justification of the homothetic transformation (20) are in order.
Physical motivation
Consider the limiting case $\lambda(x)\to\infty$, where the conformal factor approaches zero, $e^{-2\lambda(x)}\to 0$, and the associated volume form vanishes, $(d\text{vol}_{\tilde{\eta}})_{x}\to 0$. A theory that remains valid across all scales requires that the fields themselves converge to a well-defined state in this limit. We postulate that for any initial form $\alpha(x)$, the transformed field $\tilde{\alpha}(x)$ must approach a universal asymptotic form $\alpha_{d}(x)$:
@@
This principle endows the offset field $\alpha_{d}$ with a clear physical interpretation: it is the ground state to which all other forms are driven as the local spacetime metric degenerates. The existence of such a non-trivial asymptotic field is a profound consequence of the homothetic framework.
Mathematical formalism
This physical intuition can be framed rigorously using the language of contraction mappings on the cotangent bundle. At each point $x\in M$, the homothetic transformation (20) defines a fiberwise map:
@@
We endow each fiber $T_{x}^{*}M$ with the norm $\|\cdot\|_{x}$ induced by the background Minkowski metric $\eta$. For any two forms $\alpha_{1},\alpha_{2}\in T_{x}^{*}M$, the distance between their images is:
$\|\mathbf{T}_{\lambda(x)}(\alpha_{1})-\mathbf{T}_{\lambda(x)}(\alpha_{2})\|_{x}=\|e^{-\lambda(x)}(\alpha_{1}-\alpha_{2})\|_{x}=e^{-\lambda(x)}\|\alpha_{1}-\alpha_{2}\|_{x}.$
For any $\lambda(x)>0$, the Lipschitz constant $e^{-\lambda(x)}$ is strictly less than 1, meaning $\mathbf{T}_{\lambda(x)}$ is a strict contraction on the complete metric space $(T_{x}^{*}M,\|\cdot\|_{x})$. By the Banach Fixed-Point Theorem, this map has a unique fixed point. Solving for the fixed point $\alpha_{fp}$ gives:
$\alpha_{fp}=e^{-\lambda(x)}\alpha_{fp}+(1-e^{-\lambda(x)})\alpha_{d}\quad\implies\quad\alpha_{fp}=\alpha_{d}.$
Thus, for any positive scaling $\lambda(x)$, the offset form $\alpha_{d}$ is the unique, stable equilibrium point of the transformation in each fiber.
This construction defines a smooth family of fiberwise contractions across the manifold $M$, forming a pointwise contraction semigroup. As the scaling becomes infinite, the map converges to a projection that sends every form to this unique fixed point:
$\lim_{\lambda(x)\to\infty}\mathbf{T}_{\lambda(x)}(\alpha)=\alpha_{d},\quad\forall\alpha\in T_{x}^{*}M.$
This provides a rigorous mathematical justification for the physical principle that the theory must remain well-defined in the limit of a vanishing local metric. The implications of this contraction become central to our generalized Hodge theory.
How these transformations propagate through the Hodge-de Rham complex is the question that will be addressed in the next subsection, after developing some operators we need for their analysis.